HUBBLE SPACE TELESCOPE DAILY REPORT #2150 PERIOD COVERED: 0000Z (UTC) 06/22/98 - 0000Z (UTC) 06/23/98 Daily Status Report as of 174/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Three Sets of NIC/1/2/3 7817 (A Complete NICMOS Map of the Hubble Deep Field) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations to map the complete WF/PC-2 region of the central Hubble Deep Field. These data will provide unique information about the faintest objects, but will include only a small number of objects with spectroscopic redshifts or z ~ 3 Lyman break galaxies. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed Three Sets of WF/PC-2 8013 (WF/PC-2 Parallel Images to the HDF Flanking Fields {Companion Proposal to 7817}) The WF/PC-2 was used to make observations parallel to the Hubble Deep Field observations (i.e., flanking fields). The complete WF/PC-2 region of the central Hubble Deep Field {HDF} was being mapped with the NICMOS Camera 3 at F110W and F160W {proposal 7817}. While these observations were executing, the WF/PC-2 was working in parallel mode to image areas near this field. The observations were completed as scheduled, and no anomalies were reported. 1.3 Completed STIS/CCD 7926 (CCD Dark and Bias Monitor -- Continued) The Space Telescope Imaging Spectrograph (CCD) was used to make several dark and bias calibration observations. This is to measure CCD dark current and bias rates for the two supported amplifier settings {gain = 1 and gain = 4} in order to gather data for making superdarks and superbiases. These observations will also track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled, and no anomalies were noted. 1.4 Completed WF/PC-2 7620 (WF/PC-2 Cycle 7 Standard Darks) The WF/PC-2 was used to make a series of Cycle 7 standard dark calibration images. This dark calibration program obtains dark frames every week in order to provide data for the ongoing calibration of the CCD dark current rate, and to monitor and characterize the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.5 Completed Four Sets of STIS/CCD 7908 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal) The Space Telescope Imaging Spectrograph (CCD) was used to make calibration observations as part of the POMS Test Proposal. This was a STIS non-scripted parallel proposal. The observations were completed as planned, and no problems were reported. 1.6 Completed Two Sets of WF/PC-2 7909 (POMS Test Proposal: WF II Parallel Archive Proposal) The WF/PC-2 was used to make parallel observations as part of the POMS test proposal. This test proposal is designed to simulate scientific plans. The observations were completed as planned, and no problems were reported. 1.7 Completed NIC/3 7816 (NIC/3 Photometric Zero Point) The Near Infrared Camera (Camera Three) was used to make observations of standard star CSKD12-NEW in order to study the zero point flux for a set of filters. The NIC/3 intra-pixel sensitivity will also be measured. The observations were executed as scheduled, and no anomalies were noted. 1.8 Completed Two Sets of NIC/3 7919 (NICMOS Camera 3 Campaign Snapshots) The Near Infrared Camera (Camera Three) was used to make images of galaxies NGC-6808 and NGC-3782 as part of the NICMOS Camera 3 campaign snapshots. The proposal is to obtain wide field snapshot images of nearby galaxies in Paschen Alpha which is inaccessible from the ground. The observations were executed as planned, and no anomalies were reported. 1.9 Completed NIC/3 7830 (IR Spectra for Known-Mass M Dwarfs) The Near Infrared Camera (Camera Three) was used to make infrared spectrographic observations of M-Class dwarf star GJ623AB. Over the last decade there has been considerable debate on two major issues in the low-mass star literature: {1} the effective temperature scale for M dwarfs and its relation to mass and luminosity and {2} the conversion from an absolute magnitude in some passband to a bolometric magnitude. The observations were executed as scheduled, and no anomalies were noted. 1.10 Completed NIC/1/2/3 7474 (Searching for Low Mass Stars: the Mass Function at the Hydrogen Burning Limit) The Near Infrared Camera (Cameras One, Two, and Three) was used to use the unprecedented IR imaging capabilities of NICMOS to study the low-mass stars at the end of the main sequence in the nearby globular cluster NGC-6397 spanning a wide range of metallicity, precisely at the wavelength corresponding to their maximum emission. These observations would complete our ongoing investigation by making it possible, for the first time, to reliably count stars right down to and beyond the main sequence H-burning limit, thus leaving no doubt as to whether the IMF indeed drops there or starts rising once again close to the limit. The observation completed as planned with no problems. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 8 Successful : 8 Scheduled Reacquisitions : 8 Successful : 8 2.2 FHST Updates: Scheduled : 14 Successful : 14 2.3 Operations Notes: Using ROP NS-1, the SSR EDAC error counter was cleared two times. The STIS MCE-1 reset at 173/161620Z while in a level 4 SAA event and while the low voltage was off. HSTAR 6556 was written. ROP NS-12 was utilized at 173/162233Z to clear the STIC error counters. MAMA-1 will be recovered at 14824Z, the time of the next MAMA-1 high voltage turn-on. An extra DF-224 dump was collected on CCS strings A and B at 174/0636Z in accordance with Ops Note 446. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH